From the top of a 250-ft lighthouse, the angle of depression to a ship in the ocean is 28°. How far is the ship from the base of the lighthouse? Round your answer to the nearest foot.

Answer by Theo(13342)

angle of depression is 28 degrees.
90 - 28 = 62 degrees.
you have a triangle with a height of 250 perpendicular to the ground with a top angle of 62 degrees.
to find the base, find the tangent of 62 degrees.
tan(62) = opposite / adjacent = base / 250
solve for base to get:
base = 250 * tan(62) = 470.1816263
a picture of your triangle is shown below:
$$$